### v1

@@ IV-A ADMM for (5)
Let $\bm{\Lambda}\succeq 0$ and $\bm{\Theta}\succeq 0$ denote the Lagrange multipliers associated with the equality constraints in (37). Then we construct the augmented Lagrangian [52]
@@ IV-B Convergence analysis
where $\mathbf{Y}$ is the Lagrange multiplier for $\mathbf{U}\succeq 0$. It follows from (57d) that $\mathbf{Y}^{k+1}=\bm{\Lambda}^{k}-\rho(\mathbf{L}^{k+1}-\mathbf{U}^{k+1})$ must also converge to some $\mathbf{Y}^{*}$. In view of (56), we immediately have $\bm{\Lambda}^{*}=\mathbf{Y}^{*}$. Now we let $k\to\infty$, and the relations in (57) can be rewritten as
$\mathbf{L}^{*}$ $\succeq 0$ (58a)
$\bm{\Lambda}^{*}$ $\succeq 0$ (58b)
$\operatorname{tr}(\bm{\Lambda}^{*}\mathbf{L}^{*})$ $=0.$ (58c)
Similarly, a limit argument on (40f) yields $\mathbf{S}^{*}=\mathbf{V}^{*}$, and a KKT analysis on (40d) gives the conditions
$\mathbf{S}^{*}$ $\succeq 0$ (59a)
$\bm{\Theta}^{*}$ $\succeq 0$ (59b)
$\operatorname{tr}(\bm{\Theta}^{*}\mathbf{S}^{*})$ $=0.$ (59c)
@@ V-A Synthetic data examples
    Randomly generate a matrix $\bm{\Gamma}\in\mathbb{R}^{p\times r}$ with rank $r$ and a diagonal $\hat{\mathbf{S}}\succ 0$ of size $p$, where the cross-sectional dimension $p=40$.
@@ V-A1 Evaluation metrics
  • and the subspace recovery accuracy
@@ V-A2 Parameters setting
We set the tolerance parameter in (50) to be $\texttt{tol}=10^{-3}$ and the maximum number of iterations $\texttt{iter}=10^{4}$. The tuning parameters $\mu$ and $C$ represent the trade-off between the low-rankness of $\mathbf{L}^{*}$, the model misfit as measured by the KL divergence, and the sparsity of $\mathbf{S}^{*}$, while $\rho$ is the penalty parameter in the augmented Lagrangian (39). We use a CV procedure to select them. The candidate set for $C$ and $\mu$ is $\left\{10,35,60,\cdots,210\right\}$ and $\rho$ is selected from $\left\{2^{0},2^{1},\cdots,2^{5}\right\}$. We randomly partition the samples $\mathbf{y}_{i}$ generated in Step ii) above into training and validation sets of equal size. Using the optimal solution $(\tilde{\mathbf{L}},\tilde{\mathbf{S}})$ computed on the training set under a specific parameter configuration $(C,\mu,\rho)$ and the sample covariance matrix $\check{\bm{\Sigma}}^{\mathrm{v}}$ computed on the validation set, we define the following score function
@@
Next, we consider the selection of the parameter $\gamma$ in the update of $\mathbf{S}$ in (44) which comes from the proximal operator (15) and can be interpreted as a step size. We must emphasize that in our implementation, once $\gamma$ is selected, it does not change in each simulation, including the CV and the final solution with the chosen parameters $(C,\mu,\rho)$. If $\gamma$ is large, it will immediately set $\mathbf{S}$ into a zero matrix due to the thresholding operation, thus leading to convergence of the $\mathbf{S}$ update. On the other hand, when $\gamma$ is extremely small, the $\ell_{0}$ proximal operator (15) essentially reduces to an identical mapping, and Algorithm 1 is simply the ADMM for (5) without the sparsity enforcing term $C\|\mathbf{S}\|_{0}$. As a consequence, such a solution does not induce sparsity on $\mathbf{S}$. This point can also be observed from our simulations. Indeed, we take $\gamma=10^{-2},10^{-4},10^{-6},10^{-8}$, respectively, and perform Monte Carlo simulations on the same dataset with $N=1200$ and the true rank $r=10$. To speed up the convergence of Algorithm 1, a projection step onto the cone of semidefinite matrices was incorporated into the updates of $\mathbf{L}$. Fig. 2 shows the results in terms of three metrics: the KL divergence between the computed $\mathbf{L}^{*}+\mathbf{S}^{*}$ and the true covariance matrix $\bm{\Sigma}$, $\|\mathbf{S}^{*}\|_{0}$, and the running time. The subfigure 2(ii) indicates that when $\gamma=10^{-2}$, the returned $\mathbf{S}^{*}$ is a zero matrix. In contrast, when $\gamma=10^{-6}$ or $\gamma=10^{-8}$, $\mathbf{S}^{*}$ has poor sparsity. The choice $\gamma=10^{-4}$ appears to produce the best estimate $\mathbf{S}^{*}$. In addition, Table I presents the RMSE (61) of the numerical rank estimate (60) for two Monte Carlo simulations in which $N=1200$ and the true rank is $r=5$ and $r=10$, respectively. In view of the results in Fig. 2 and Table I, we set $\gamma=10^{-4}$ which seems rather reasonable for synthetic datasets.
@@
Figure 2: The performance of the $\ell_{0}$ ADMM in terms of three metrics when $\gamma$ takes the value of $10^{-2}$, $10^{-4}$, $10^{-6}$, and $10^{-8}$, respectively, and $N=1200$, the true $r=10$.
V-A3 Comparison with [15]
Similar to the simulation setup in [15], we consider two scenarios in which the true rank of $\bm{\Gamma}$ is $r=5$ and $r=10$, respectively, and perform Monte Carlo simulations with different data lengths $N$. The subspace recovery index (62) for each trial is shown via the boxplot in Fig. 3. One can observe that the subspace recovery accuracy increases with the number of available samples. Moreover, despite the presence of a few outliers, the values of $\mathrm{ratio}(\mathbf{\Gamma}^{*})$ consistently surpass the $0.995$ threshold. Compared to the prior work [15], more precisely, Fig. 2 in that paper, we remark that Algorithm 1 has made visible improvements in the subspace recovery of $\bm{\Gamma}$.
In addition, we have also generated a set of samples of size $N=1200$ and $r=5$ to see the effect of parameters $C$, $\mu$, and $\rho$ on the subspace recovery index. More specifically, Fig. 4 illustrates the quantity $\mathrm{ratio}(\mathbf{\Gamma}^{*})$ as a function of $(C,\mu)$ for a fixed $\rho=16$. The surface plot is quite flat which means that Algorithm 1 achieves robust recovery of the column space of $\bm{\Gamma}$ under different parameter configurations $(C,\mu)$ which are used in the CV. We add that the figure basically does not change for different values of $\rho$ (with the same range for $(C,\mu)$) and hence they are not reported here.
V-A4 Comparison with the $\ell_{1}$ ADMM
@@
and the matrix form is also an elementwise application of the scalar form above. The $\ell_{1}$ proximal operator is also called soft thresholding operator in the literature.
@@
When the true $\hat{\mathbf{S}}$ is set as the identity matrix, panels (a) and (b) of Fig. 5 show graphically the estimates $\mathbf{S}^{*}$ computed from the $\ell_{0}$ and the $\ell_{1}$ versions of the ADMM with $\gamma=10^{-4}$ fixed. We note that the $\ell_{0}$ ADMM correctly recognizes the underlying diagonal structure of $\hat{\mathbf{S}}$, while the $\ell_{1}$ version returns a zero matrix which is similar to the outcome of the $\ell_{0}$ ADMM for $\gamma=10^{-2}$. See the comment in Subsubsec. V-A2. Panels (c) and (d) of Fig. 5 show the results of the $\ell_{1}$ ADMM corresponding to $\gamma=10^{-6}$ and $\gamma=10^{-8}$, respectively. When $\gamma=10^{-6}$, the $\ell_{1}$ ADMM returns only a few nonzero entries along the diagonal of $\mathbf{S}^{*}$ with small numerical values and fails to reveal the diagonal structure of $\hat{\mathbf{S}}$. For the case of $\gamma=10^{-8}$, the $\ell_{1}$ ADMM treats the optimization problem that essentially neglects the sparsity enforcing term $C\|\mathbf{S}\|_{1}$, much similar to the $\ell_{0}$ case. Indeed, both the $\ell_{0}$ and the $\ell_{1}$ proximal operators reduce to the identity mapping when $\gamma\to 0$. The simulation results are also not far from each other: the $\ell_{1}$ estimate of $\mathbf{S}^{*}$ has many spurious off-diagonal nonzeros (see panel (d) of Fig. 5) and the situation of the $\ell_{0}$ estimate can be seen from panel (ii) of Fig. 2. In summary, the $\ell_{0}$ ADMM with a suitable choice of $\gamma$ appears to be able to identify the sparse structure of $\hat{\mathbf{S}}$ in a more reliable fashion.
(a) $r=5$
@@
Figure 3: The subspace recovery accuracy of $\bm{\Gamma}$ for $N=400,800,1200$ in two scenarios: the true rank $r=5$ and $r=10$, respectively.
Figure 4: The subspace recovery accuracy across various parameter configurations $(\mu,C)$ with $\rho=16$ fixed on a dataset of size $N=1200$ and the true $r=10$
Figure 5: Comparison of the sparse structure of the estimated $\mathbf{S}^{*}$ under the $\ell_{0}$ and $\ell_{1}$ regularizations, respectively, when $N=1200$, the true $\hat{\mathbf{S}}=\mathbf{I}$ and $r=5$. (a) $\ell_{0}$ regularization, $\gamma=10^{-4}$, (b, c, d) $\ell_{1}$ regularization, $\gamma=10^{-4},10^{-6}$, and $10^{-8}$, respectively.

### v2

@@ IV-A ADMM for (5)
Let $\bm{\Lambda}$ and $\bm{\Theta}$ denote the Lagrange multipliers associated with the equality constraints in (37). Then we construct the augmented Lagrangian [52]
@@ IV-B Convergence analysis
where $\mathbf{Y}$ is the Lagrange multiplier for $\mathbf{U}\succeq 0$. It follows from (57d) that $\mathbf{Y}^{k+1}=\bm{\Lambda}^{k}-\rho(\mathbf{L}^{k+1}-\mathbf{U}^{k+1})$ must also converge to some $\mathbf{Y}^{*}$. In view of (56), we immediately have $\bm{\Lambda}^{*}=\mathbf{Y}^{*}$. Now we let $k\to\infty$, and the relations in (57) can be rewritten as $\mathbf{L}^{*}\succeq 0$, $\bm{\Lambda}^{*}\succeq 0$, and $\operatorname{tr}(\bm{\Lambda}^{*}\mathbf{L}^{*})=0$.
Similarly, a limit argument on (40f) yields $\mathbf{S}^{*}=\mathbf{V}^{*}$, and a KKT analysis on (40d) gives the conditions $\mathbf{S}^{*}\succeq 0$, $\bm{\Theta}^{*}\succeq 0$, and $\operatorname{tr}(\bm{\Theta}^{*}\mathbf{S}^{*})=0$.
@@ V-A Synthetic data examples
    Randomly generate a matrix $\bm{\Gamma}\in\mathbb{R}^{p\times r}$ with rank $r$ and a sparse $\hat{\mathbf{S}}\succ 0$ of size $p$, where the cross-sectional dimension $p=40$.
@@ V-A1 Evaluation metrics
  • the subspace recovery accuracy
@@ V-A2 Parameters setting
We set the tolerance parameter in (50) to be $\texttt{tol}=10^{-3}$ and the maximum number of iterations $\texttt{iter}=10^{4}$. The tuning parameters $\mu$ and $C$ represent the trade-off between the low-rankness of $\mathbf{L}^{*}$, the model misfit as measured by the KL divergence, and the sparsity of $\mathbf{S}^{*}$, while $\rho$ is the penalty parameter in the augmented Lagrangian (39). We use a CV procedure to select them. The candidate set for $C$ and $\mu$ is $\left\{60,110,\cdots,360\right\}$ and $\rho$ is selected from $\left\{2^{0},2^{1},\cdots,2^{5}\right\}$. We randomly partition the samples $\mathbf{y}_{i}$ generated in Step ii) above into training and validation sets of equal size. Using the optimal solution $(\tilde{\mathbf{L}},\tilde{\mathbf{S}})$ computed on the training set under a specific parameter configuration $(C,\mu,\rho)$ and the sample covariance matrix $\check{\bm{\Sigma}}^{\mathrm{v}}$ computed on the validation set, we define the following score function
@@
Next, we consider the selection of the parameter $\gamma$ in the update of $\mathbf{S}$ in (44) which comes from the proximal operator (15) and can be interpreted as a step size. We must emphasize that in our implementation, once $\gamma$ is selected, it does not change in each simulation, including the CV and the final solution with the chosen parameters $(C,\mu,\rho)$. If $\gamma$ is large, it will immediately set $\mathbf{S}$ into a zero matrix due to the thresholding operation, thus leading to convergence of the $\mathbf{S}$ update. On the other hand, when $\gamma$ is extremely small, the $\ell_{0}$ proximal operator (15) essentially reduces to an identical mapping, and Algorithm 1 is simply the ADMM for (5) without the sparsity enforcing term $C\|\mathbf{S}\|_{0}$. As a consequence, such a solution does not induce sparsity on $\mathbf{S}$. This point can also be observed from our simulations. Indeed, we take $\gamma=10^{-2},10^{-4},10^{-6},10^{-8}$, respectively, and perform Monte Carlo simulations on the same dataset with $N=1000$, the true rank $r=4$, and a signal-to-noise ratio (SNR, defined as $\|\bm{\Gamma}\bm{\Gamma}^{\top}\|_{\mathrm{F}}/\|\hat{\mathbf{S}}\|_{\mathrm{F}}$) equal to $6$. To speed up the convergence of Algorithm 1, a projection step onto the cone of semidefinite matrices was incorporated into the updates of $\mathbf{L}$. Fig. 2 shows the results in terms of three metrics: the KL divergence between the computed $\mathbf{L}^{*}+\mathbf{S}^{*}$ and the true covariance matrix $\bm{\Sigma}$, $\|\mathbf{S}^{*}\|_{0}$, and the running time. The subfigure 2(ii) indicates that when $\gamma=10^{-2}$, the returned $\mathbf{S}^{*}$ is a zero matrix. In contrast, when $\gamma=10^{-6}$ or $\gamma=10^{-8}$, $\mathbf{S}^{*}$ has poor sparsity. The choice $\gamma=10^{-4}$ appears to produce the best estimate $\mathbf{S}^{*}$. In addition, Table I presents the RMSE (59) of the numerical rank estimate (58) for two Monte Carlo simulations in which $N={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1000}$ and the true rank is $r={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}$ and $r=10$, respectively. In view of the results in Fig. 2 and Table I, the value $\gamma=10^{-4}$ seems reasonable for synthetic datasets with $\mathrm{SNR}=6$.
@@
Figure 2: The performance of the $\ell_{0}$ ADMM in terms of three metrics when $\gamma$ takes the value of $10^{-2}$, $10^{-4}$, $10^{-6}$, and $10^{-8}$, respectively, with $N=1000$, the true $r=4$, $\mathrm{SNR}=6$ and the sparsity level of $\hat{\mathbf{S}}$ being $\|\hat{\mathbf{S}}\|_{0}/p^{2}=0.055$.
V-A3 Comparison with the $\ell_{1}$ ADMM
@@
and the matrix form is also an elementwise application of the scalar form above. The $\ell_{1}$ proximal operator is also called soft thresholding operator in the literature. Such a comparison is also motivated by the fact that the globally optimal solution to the $\ell_{1}$ relaxed optimization problem is not even locally optimal in the $\ell_{0}$ sense, see [57, Theorem 1] for the precise statement.
@@
We separately consider two structures of the true $\hat{\mathbf{S}}$ under $\mathrm{SNR}=6$: diagonal and nondiagonal sparse. When $\hat{\mathbf{S}}$ is set as a diagonal matrix, more precisely, a scaled identity matrix of the form $\alpha\mathbf{I}$ with $\alpha>0$, panel (a) of Fig. 3 shows graphically the estimates $\mathbf{S}^{*}$ computed from the $\ell_{0}$ and the $\ell_{1}$ versions of the ADMM at different values of $\gamma$. In the case of $\gamma=10^{-4}$, we note that the $\ell_{0}$ ADMM correctly recognizes the underlying diagonal structure of $\hat{\mathbf{S}}$, while the $\ell_{1}$ version returns a zero matrix which is similar to the outcome of the $\ell_{0}$ ADMM for $\gamma=10^{-2}$. See the comment in Subsubsec. V-A2. When $\gamma=10^{-6}$, the $\ell_{1}$ ADMM returns only a few nonzero entries along the diagonal of $\mathbf{S}^{*}$ with small numerical values and fails to reveal the diagonal structure of $\hat{\mathbf{S}}$. For the case of $\gamma=10^{-8}$, the $\ell_{1}$ ADMM treats the optimization problem that essentially neglects the sparsity enforcing term $C\|\mathbf{S}\|_{1}$, much similar to the $\ell_{0}$ case. Indeed, both the $\ell_{0}$ and the $\ell_{1}$ proximal operators reduce to the identity mapping when $\gamma\to 0$. The simulation results are also not far from each other: the $\ell_{1}$ estimate of $\mathbf{S}^{*}$ has many spurious off-diagonal nonzeros and the situation of the $\ell_{0}$ estimate can be seen from panel (ii) of Fig. 2. Moreover, when $\hat{\mathbf{S}}$ is set as a nondiagonal sparse matrix, panel (b) of Fig. 3 clearly demonstrates the superior performance of the $\ell_{0}$ ADMM in estimating the true sparse structure. In summary, the $\ell_{0}$ ADMM with a suitable choice of $\gamma$ appears to be able to identify the sparse structure of $\hat{\mathbf{S}}$ in a more reliable fashion.
Figure 3: Comparison of the sparse structure of the estimated $\mathbf{S}^{*}$ under the $\ell_{0}$ and $\ell_{1}$ regularizations across different conditions, when $N=1000$, $r=4$ and $\mathrm{SNR}=6$. (a) The true $\hat{\mathbf{S}}$ is diagonal (scaled identity). (b) The true $\hat{\mathbf{S}}$ is nondiagonal sparse.
V-A4 Comparison with [15]
Similar to the simulation setup in [15], we set the true sparse component $\hat{\mathbf{S}}$ to be a scaled identity matrix in step i), and $\mathrm{SNR}=1$. The parameter $\gamma$ is set to $10^{-2}$ in this case. Then, we consider two scenarios in which the true rank of $\bm{\Gamma}$ is $r=4$ and $r=10$, respectively, and perform Monte Carlo simulations consisting of $200$ trials with different data lengths $N$. The subspace recovery index (60) for each trial is shown via the boxplot in Fig. 4. One can observe that the subspace recovery accuracy increases with the number of available samples. Moreover, compared to Fig. 2 in the prior work [15], panel (b) of our Fig. 4 shows that the method in [15] performs better in the subspace recovery of $\bm{\Gamma}$ when the true rank is $r=10$. However, panel (a) shows that in the case of $r=4$, the opposite happens, i.e., our Algorithm 1 achieves better performance, even when the outliers (marked with red “+”) are taken into consideration.
In addition, we have also generated a set of samples of size $N=1000$ and $r=4$ to see the effect of parameters $C$, $\mu$, and $\rho$ on the subspace recovery index. More specifically, Fig. 5 illustrates the quantity $\mathrm{ratio}(\mathbf{\Gamma}^{*})$ as a function of $(C,\mu)$ for a fixed $\rho=16$. The surface plot is quite flat which means that Algorithm 1 achieves robust recovery of the column space of $\bm{\Gamma}$ under different parameter configurations $(C,\mu)$ which are used in the CV. We add that the figure basically does not change for different values of $\rho$ (with the same range for $(C,\mu)$) and hence they are not reported here.
(a) $r=4$
@@ V-A4 Comparison with [15]
Figure 4: The subspace recovery accuracy of $\bm{\Gamma}$ for $N=200$, $500$, and $1000$ in two scenarios: when the true rank $r=4$ and when $r=10$, respectively, with the true $\hat{\mathbf{S}}$ being diagonal (scaled identity) and $\mathrm{SNR}=1$. The integer below each boxplot represents the total number of outliers (marked with red “+”).
Figure 5: The subspace recovery accuracy across various parameter configurations $(\mu,C)$ with $\rho=16$ fixed on a dataset of size $N=1000$, the true $r=4$ and $\mathrm{SNR}=1$.
@@ VI Conclusion
For future work, it could be interesting to compare our approach with reweighted $\ell_{1}$ minimization [60] and other optimization models involving nonconvex relaxations of the $\ell_{0}$ norm.